# Confinement-Higgs and deconfinement-Higgs transitions in three-dimensional $Z(2)$ LGT

### v1

@@ 1 Introduction
Two order parameters can distinguish deconfinement from confinement and Higgs phases. The first of these order parameters is the Fredenhagen-Marcu (FM) operator introduced in Ref.[10] and tested in Monte Carlo simulations in [11] on the $4d$ $Z(2)$ gauge-Higgs theory, which has a similar phase structure. The rigorous proof that the FM operator behaves in the expected way was provided recently in Ref.[12].
@@ 3.4 The FM operator
The observed dependence on the distance can be summarized as follows: below a certain critical value of $\gamma$, the ratio drops rapidly towards zero for increasing $R$, whereas at larger values of $\gamma$ it approaches a constant value, which increases with increasing $\gamma$. A careful inspection of the plot and the Table for $\gamma$ values below the phase transition at $\beta=0.9$, together with a fitting, reveals, however, that the FM operator becomes very small at large distances $R$ but it is nonvanishing. In order to check if this is the case, we performed simulations on several lattices sizes $L=16,32,48,64$ for three $\gamma$ values in the deconfinement phase. Those results are shown in Fig.16. These plots indicate that the small nonzero value of the FM operator is rather a finite-size effect: the FM operator tends to zero as expected for increasing lattice size. We did not study the scaling behavior in the vicinity of the critical point as it requires a huge numerical effort.
@@
2 $1.05109(26)\cdot 10^{-2}$ $1.99735(64)\cdot 10^{-2}$ $9.3183(24)\cdot 10^{-2}$
4 $6.1345(75)\cdot 10^{-4}$ $2.8599(28)\cdot 10^{-3}$ $6.2745(27)\cdot 10^{-2}$
6 $7.996(22)\cdot 10^{-5}$ $8.187(17)\cdot 10^{-4}$ $5.6525(32)\cdot 10^{-2}$
8 $2.195(12)\cdot 10^{-5}$ $3.1392(88)\cdot 10^{-4}$ $5.3988(30)\cdot 10^{-2}$
10 $1.3618(78)\cdot 10^{-5}$ $1.4620(52)\cdot 10^{-4}$ $5.2574(32)\cdot 10^{-2}$
12 $1.2205(73)\cdot 10^{-5}$ $7.987(34)\cdot 10^{-5}$ $5.1654(31)\cdot 10^{-2}$
14 $1.2067(63)\cdot 10^{-5}$ $5.050(24)\cdot 10^{-5}$ $5.1000(34)\cdot 10^{-2}$
16 $1.2195(58)\cdot 10^{-5}$ $3.584(16)\cdot 10^{-5}$ $5.0534(37)\cdot 10^{-2}$
18 $1.2343(70)\cdot 10^{-5}$ $2.706(14)\cdot 10^{-5}$ $5.0197(32)\cdot 10^{-2}$
20 $1.2284(93)\cdot 10^{-5}$ $1.983(11)\cdot 10^{-5}$ $4.9968(35)\cdot 10^{-2}$
22 $1.2430(58)\cdot 10^{-5}$ $1.213(10)\cdot 10^{-5}$ $4.9833(38)\cdot 10^{-2}$
24 $1.2665(57)\cdot 10^{-5}$ $7.754(42)\cdot 10^{-7}$ $4.9797(36)\cdot 10^{-2}$
@@
Figure 16: FM operator at $\beta=0.9$ in the deconfinement phase.
@@ 4 Summary
Acknowledgments. The authors acknowledge support from INFN/NPQCD project. This work is partially supported by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU. Most numerical simulations have been performed on the CSN4 cluster of the Scientific Computing Center at INFN-Pisa. O.B. thanks the Pisa section of INFN for the warm hospitality during a working stay. B.A. is grateful to Juan José Alonso for useful advice on simulation methods for spin-glasses. B.A. and O.B. also thank Claudio Bonati, Massimo D’Elia and Claudio Bonanno for discussions.

### v2

@@ 1 Introduction
Two order parameters can distinguish deconfinement from confinement and Higgs phases. The first of these order parameters is the Fredenhagen-Marcu (FM) operator introduced in Ref.[10] and tested in Monte Carlo simulations in [11] on the $4d$ $Z(2)$ gauge-Higgs theory, which has a similar phase structure. The rigorous proof that the FM operator behaves in the expected way was provided recently in Ref.[12]. The FM operator has been studied also in Ref. [13], in a closely related toric code model, which has a similar phase structure.
@@ 3.4 The FM operator
The observed dependence on the distance can be summarized as follows: below a certain critical value of $\gamma$, the ratio drops rapidly towards zero for increasing $R$, whereas at larger values of $\gamma$ it approaches a constant value, which increases with increasing $\gamma$. While the approach to a constant can be easily identified, the approach to zero is a more delicate problem. In order to be sure that the FM operator does vanish in the deconfinement phase, we performed simulations on several lattices sizes $L=16,32,48,64$ for three $\gamma$ values in the deconfinement phase. Results obtained are shown on Fig. 16. These plots clearly indicate that, within errors, the FM operator does vanish in the deconfinement phase as expected. We did not study the scaling behavior in the vicinity of the critical point as it requires a huge numerical effort.
@@
2 $1.04970(26)\cdot 10^{-2}$ $1.99443(63)\cdot 10^{-2}$ $4.6162(18)\cdot 10^{-2}$
4 $6.0000(75)\cdot 10^{-4}$ $2.8233(27)\cdot 10^{-3}$ $1.9426(16)\cdot 10^{-2}$
6 $6.742(20)\cdot 10^{-5}$ $7.817(16)\cdot 10^{-4}$ $1.3657(14)\cdot 10^{-2}$
8 $9.846(94)\cdot 10^{-6}$ $2.7844(96)\cdot 10^{-4}$ $1.0958(13)\cdot 10^{-2}$
10 $1.639(33)\cdot 10^{-6}$ $1.1288(52)\cdot 10^{-4}$ $9.185(12)\cdot 10^{-3}$
12 $2.79(15)\cdot 10^{-7}$ $4.897(34)\cdot 10^{-5}$ $7.800(11)\cdot 10^{-3}$
14 $4.34(50)\cdot 10^{-8}$ $2.204(50)\cdot 10^{-5}$ $6.619(11)\cdot 10^{-3}$
16 $7.5(2.5)\cdot 10^{-9}$ $9.91(16)\cdot 10^{-6}$ $5.5755(91)\cdot 10^{-3}$
18 $1.9(1.2)\cdot 10^{-9}$ $4.34(10)\cdot 10^{-6}$ $4.6412(80)\cdot 10^{-3}$
20 -- $1.553(48)\cdot 10^{-6}$ $3.8068(69)\cdot 10^{-3}$
22 -- $3.46(16)\cdot 10^{-7}$ $3.0654(58)\cdot 10^{-3}$
24 -- -- $2.4151(46)\cdot 10^{-3}$
@@
Figure 16: FM operator at $\beta=0.9$ in the deconfinement phase. Data at different $L$ have been slightly shifted along the horizontal axix for clarity purposes.
@@ 4 Summary
Acknowledgments. The authors acknowledge support from INFN/NPQCD project. This work is partially supported by ICSC - National Centre for HPC, Big Data and Quantum Computing, funded by European Union – NextGenerationEU. Most numerical simulations have been performed on the CSN4 cluster of the Scientific Computing Center at INFN-Pisa. O.B. thanks the Pisa section of INFN for the warm hospitality during a working stay. B.A. is grateful to Juan José Alonso for useful advice on simulation methods for spin-glasses. B.A. and O.B. also thank Claudio Bonati, Massimo D’Elia and Claudio Bonanno for discussions.